OpenStudy (anonymous):

Find cos θ if sin θ = -5/13 and tan θ > 0.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@sidsiddhartha

hartnn (hartnn):

\(\cos^2 \theta = 1- \sin^2 \theta \)

OpenStudy (anonymous):

i got -13/5

hartnn (hartnn):

how ?

OpenStudy (anonymous):

for the equation

hartnn (hartnn):

\(\cos \theta = \sqrt{1-(-5/13)^2} \)

hartnn (hartnn):

\(\cos \theta = \sqrt {144/169} = 12/13\)

hartnn (hartnn):

and since tan theta is positive both sin and cos would be negative so final answer would be cos theta = -12/13

OpenStudy (anonymous):

ohh!!!